OpenStudy (anonymous):

use the quadratic formula to find the zero of f(x)=x^2-x-6

OpenStudy (anonymous):

a=1 b=-1 c=-6 \[(-b \pm \sqrt{b ^{2}-4ac})/(2a)\] \[(1\pm \sqrt{(-1)^{2}-4(1)(-6)})/(2)(1)\] \[(1\pm \sqrt{1+24})/2\] \[(1\pm5)/2\] 6/2=3, -4/2=-2 3,-2